# Inv-Entropy: A Fully Probabilistic Framework for Uncertainty Quantification in Language Models

### v1

@@ 3 Experiments
Our framework requires three inputs: $\psi$, $\mathcal{P}er$, and $a_{\text{Similarity}}$. For $\mathcal{P}er$, we apply two strategies: (1) ChatGPT-based paraphrasing, generating nine perturbed versions per question, (2) GAAP introduced in Sec. 2.5 with a similarity threshold of $\delta=0.7$. For $\psi$, we employ three state-of-the-art approaches: (i) SBERT-small (paraphrase-MiniLM-L6-v2), (ii) SBERT-large (all-mpnet-base-v2) (Reimers and Gurevych, 2019), and (iii) DeBERTa-v2-xlarge-MNLI (He et al., 2021). For (i) and (ii), we use cosine similarity $a_{\text{Similarity}}(x,x^{\prime})=(1+\cos(x,x^{\prime}))/2$. While, (iii) generates an entailment score; however, this score is not symmetric. To address this, we take the average $(a_{\text{Similarity}}(x,x^{\prime})+a_{\text{Similarity}}(x^{\prime},x))/2$. All experiments were conducted with an NVIDIA A100 GPU. Detailed experimental set up including prompts and parameters used are detailed in Appendix D. Appendix D also includes additional simulation results; we present only the core findings in the main paper for clarity and focus.
@@ 3.1 Results
Calibration-based
As shown in Table 1, our proposed Inv-Entropy method demonstrates consistently strong and stable performance across all five datasets and metrics. While it does not always rank first, it frequently appears among the top performers and achieves notable results, such as 0.788 AUROC on TriviaQA and 0.898 PRR on MMLU. It also ranks among the top two in Brier score, indicating well-calibrated confidence estimates. Its variant, NR-Inv-Entropy, also performs competitively, underscoring the robustness of our approach and the added benefit of incorporating replications. We intentionally use ChatGPT-based paraphrasing to highlight the advantages of our method independent of GAAP. Results for LLaMA-3.1-8B-Instruct are deferred to Appendix D.
@@
Table 2 reports TSU results across all five datasets. These results align with the calibration-based UQ metrics in Table 1: methods defined by our probabilistic framework consistently achieve top performance, while LexSim, DegMat and LUQ exhibit highly inconsistent results, occasionally leading in specific cases but underperforming in others. Methods like P(True) perform poorly in TSU due to their binary nature, which fundamentally lacks granularity. Similarly, Semantic Entropy shows limited discriminatory power as its values often cluster around few discrete points. In contrast to these coarse-grained approaches, INV-Entropy is fine-grained with superior ability to reflect uncertainty variations, demonstrated by consistent top TSU performance across all datasets and temperature settings. This result highlights the probabilistic framework’s ability to capture uncertainty trends beyond labeled datasets.
@@
(i) Impact of replications: The comparison between NR-Inv-Entropy and Inv-Entropy in Tables 1 and 2 highlights the impact of replication and bootstrapping. Notably, NR-Inv-Entropy often ranks second, showcasing the strength of our framework even without replication. This suggests that one can trade off a small loss in accuracy for fewer queries. Appendix D includes an ablation over $S$ and $r$, further confirming this finding. (ii) Impact of inversion: The comparison between NI-Entropy and Inv-Entropy further confirms the advantages of our inverse approach, which examines the diversity of inputs that could have led to a specific output. (iii) Framework generality: The often competitive performance of WD-px-py and MAX-py-x highlights the robustness of our framework and its flexibility in defining a wide range of UQ measures. (iv) Impact of perturbations: Fig. 4 shows that GAAP consistently improves AUROC across all three datasets and both LLMs compared to ChatGPT-based paraphrasing. By principledly enhancing input diversity, GAAP better tests model robustness and improves uncertainty quantification. These results underscore the importance of meaningful perturbations. (v) Impact of embedding function: Our framework delivers consistently strong uncertainty estimates with every encoder we tested, including SBERT (paraphrase-MiniLM-L6-v2), SBERT (all-mpnet-base-v2), and DeBERTa. Although the larger encoders provide slightly higher scores on several tasks, the overall gap is modest, showing that even lightweight models can support reliable UQ when paired with our method. The close alignment between results from SBERT (all-mpnet-base-v2) and DeBERTa further suggests that entailment and similarity signals extracted by the two architectures contain overlapping information.
@@
[…]
Method
TSU (0.3, 0.7)
TSU (0.7, 1.0)
TSU (1.0, 1.4)
TSU (0.3, 1.0)
TSU (0.3, 1.4)
TSU (0.7, 1.4)
TSU (0.3-1.0)
TSU (0.7-1.4)
TSU (0.3-1.4)
TriviaQA
Semantic Entropy 11.57 14.26 17.35 21.56 25.51 17.35 5.18 5.35 3.94
VU 22.45 27.55 38.78 25.51 41.84 42.86 0.00 4.92 0.00
P(True) 1.02 2.13 3.85 1.38 0.98 3.49 0.00 0.00 0.00
LexSim 22.39 22.95 46.94 30.61 53.36 53.54 9.18 12.38 8.16
DegMat 24.58 33.21 45.37 31.77 48.98 47.96 18.37 20.02 13.27
LUQ 20.41 31.08 48.06 33.67 52.34 50.00 14.78 27.63 10.20
KLE 7.14 17.35 13.45 4.57 1.28 6.42 2.79 1.31 0.00
Inv-Entropy 52.42 66.33 77.55 76.53 92.86 88.78 30.49 47.21 19.05
NI-Entropy 55.14 47.96 61.22 59.35 67.35 60.20 21.43 19.32 11.73
WD-px-py 44.81 67.35 57.62 60.20 64.49 69.39 23.11 36.73 12.06
MAX-py-x 65.31 50.83 74.49 67.35 85.71 81.63 27.42 32.65 16.03
SciQ
Semantic Entropy 11.09 15.48 21.55 17.35 27.62 25.51 6.73 4.31 2.65
VU 21.43 33.67 28.57 35.71 31.42 32.65 1.55 6.12 0.00
P(True) 0.00 3.42 3.14 0.00 1.33 0.79 0.00 0.00 0.00
LexSim 31.73 29.59 47.96 40.82 61.22 59.18 19.39 13.27 11.22
DegMat 39.80 47.96 57.14 52.04 64.28 64.46 23.47 23.99 13.04
LUQ 38.45 43.88 41.84 44.90 57.14 44.95 18.37 17.35 6.12
KLE 6.12 39.54 19.32 4.98 5.67 18.37 2.30 5.10 0.00
Inv-Entropy 46.70 62.24 72.38 66.72 75.84 80.61 22.45 35.71 14.31
NI-Entropy 42.30 59.18 55.19 48.98 54.77 63.27 21.43 22.65 9.18
WD-px-py 50.00 59.16 63.25 59.37 71.40 69.39 22.13 29.59 11.22
MAX-py-x 51.02 61.22 67.35 61.22 69.86 73.25 29.59 34.69 15.29
NQ
Semantic Entropy 22.53 20.41 35.71 27.55 39.80 30.78 4.08 10.20 2.04
VU 23.71 22.68 40.21 21.65 36.08 36.08 4.12 3.09 0.00
P(True) 4.08 3.06 2.04 1.02 2.04 1.02 0.00 0.00 0.00
LexSim 51.55 54.64 59.79 63.73 79.38 74.23 36.08 32.99 17.44
DegMat 51.04 51.04 59.38 65.62 76.04 71.88 32.29 33.33 21.42
LUQ 43.16 49.47 67.37 57.89 76.84 74.74 24.21 30.41 13.68
KLE 6.38 28.72 21.28 6.38 4.26 12.77 2.13 4.26 0.00
Inv-Entropy 61.86 59.79 78.35 74.23 91.75 85.57 30.93 41.24 22.43
NI-Entropy 47.42 60.82 59.79 58.76 70.10 70.10 23.71 29.81 10.31
WD-px-py 63.55 50.52 70.10 67.01 76.29 72.16 27.84 25.77 16.92
MAX-py-x 55.67 62.89 73.20 66.30 84.54 82.47 28.87 39.18 18.56
MMLU
Semantic Entropy 21.43 17.35 33.20 21.83 42.08 39.80 7.14 4.93 2.09
VU 23.56 19.01 37.62 30.18 32.50 38.73 1.37 2.59 0.00
P(True) 1.92 4.56 5.87 4.92 5.02 5.79 0.00 0.00 0.00
LexSim 33.94 41.17 55.06 50.00 68.37 61.22 24.78 30.61 15.28
DegMat 41.84 54.08 69.39 53.76 78.57 77.55 21.46 32.58 14.34
LUQ 53.46 47.96 61.22 58.92 68.37 62.24 27.55 27.55 10.80
KLE 10.20 25.51 26.53 7.14 4.76 12.23 2.93 2.67 0.00
Inv-Entropy 60.08 67.35 73.47 79.59 90.82 86.73 34.31 43.88 18.37
NI-Entropy 56.12 54.63 50.00 67.35 52.45 59.38 19.39 18.37 7.14
WD-px-py 50.00 59.16 66.38 59.37 70.41 69.39 25.60 29.59 11.22
MAX-py-x 62.24 62.35 65.41 73.47 81.73 80.61 31.62 25.51 14.42
GSM8K
Semantic Entropy 44.90 56.12 35.71 71.43 77.55 62.24 20.41 13.27 4.08
VU 11.34 39.18 29.90 35.05 35.05 38.14 2.06 6.19 1.03
P(True) 5.10 17.35 6.12 3.32 3.98 11.22 0.00 0.00 0.00
LexSim 54.17 63.54 54.17 65.62 64.58 63.54 30.31 23.96 10.42
DegMat 55.79 64.21 55.79 72.63 75.79 70.53 29.47 29.47 10.53
LUQ 64.95 74.23 72.16 83.51 89.69 81.44 44.33 54.64 37.11
KLE 17.02 35.11 27.66 4.23 3.72 23.40 0.00 3.19 0.00
Inv-Entropy 73.68 67.33 68.42 86.32 93.68 72.63 45.81 44.21 30.53
NI-Entropy 53.61 58.76 51.55 61.16 57.73 59.79 25.77 20.62 6.19
WD-px-py 54.08 53.06 68.37 55.10 64.29 60.20 20.41 27.43 7.14
MAX-py-x 71.88 65.62 62.50 89.58 91.67 77.08 42.71 37.50 28.12
@@ Appendix A Related Work
UQ in LLMs has been gaining increasing interest recently [Huang et al., 2024]. The existing, albeit limited, literature can generally be categorized into three perspectives.
@@
Self-evaluation techniques for LLMs [Chen et al., 2024] and the verbal expression of uncertainty [Lin et al., 2022] have recently been explored to enhance interpretability and reliability. For instance, recent approaches for evaluating free-form generation tasks frequently utilize LLMs as evaluators [Zheng et al., 2023]. Additionally, for gray-box models, where the internal workings are partially known, perplexity [Chen et al., 1998] and entropy [Shannon, 1948] can be directly computed from the output logits, providing a natural UQ measure.
@@
These methods generate multiple outputs for a given input and measure the deviation between them to estimate uncertainty [Grewal et al., 2024, Wagner et al., 2024, Kuhn et al., 2023]. Perhaps most prevalent is semantic entropy [Farquhar et al., 2024], which computes uncertainty by clustering semantically equivalent answers from multiple responses and calculating the entropy of the resulting clusters. While effective in capturing aleatoric uncertainty, these approaches struggle with confidently wrong predictions, as resampling often yields similar incorrect results, leading to overconfidence and poor calibration. This issue exacerbates the challenges of handling hallucinations in LLMs [Grewal et al., 2024].
@@
This is a more recent approach that involves systematic perturbations of inputs or latent representations to evaluate output variability [Zhang et al., 2023]. While Dong et al. [2023] systematically evaluates LLM robustness for noisy slot-filling tasks under diverse input perturbations, SPUQ [Gao et al., 2024] provides a UQ metric by analyzing response variations to perturbed inputs. Notably, SPUQ achieves significant improvements in model uncertainty calibration, reducing Expected Calibration Error (ECE) by an average of 50%. Indeed, our theoretical argument in Sec. 2.1 provides a theoretical justification for the need for perturbation in effective UQ.
@@
By assumption, there exists a constant $G>0$ such that $\lVert\nabla^{2}f(x)\rVert_{op}\leq G$, $\forall x\in\mathbb{R}^{d}$. Therefore,
@@
Since $z$ admits multivariate Gaussian distribution, we can provide an upper bound for $\mathbb{E}(R(z))$ and $\mathbb{E}(R^{2}(z))$ as
@@
$=\mathbb{E}\left[\left(\nabla\hat{f}(x_{0})^{\top}(I-\eta_{\nabla f^{\star}})z\right)^{2}\right]+\mathbb{E}\left[\left(\nabla\hat{f}(x_{0})^{\top}(I-\eta_{\nabla f^{\star}})z\right)\zeta(z)\right]+\mathbb{E}\left[\zeta(z)^{2}\right],$
@@ C.1 An example
Table 3: A question $S_{x_{0}}$ and $10$ perturbed versions of $S_{x_{0}}$ as the output of GAAP.
@@ C.2.1 Key words selection
where the indices $\{j_{1},j_{2},\dots,j_{q}\}\subseteq\{1,\dots,p\}$, and the number of selected key tokens are: $q=\text{int}[r\cdot p]$. In GAAP, we use KeyBERT [Grootendorst, 2024] to implement $k(\cdot,\cdot)$.
@@ C.2.3 Iterative population update
    In the terminology of genetic algorithms, we define our fitness function as the semantic similarity to $x_{0}$, $a_{\text{Similarity}}(x_{0},\cdot)$. Then, we compute the fitness value $a_{\text{Similarity}}(x_{0},x_{i})$ for $\forall S_{x_{i}}\in\text{Pop}_{t}(S_{x_{0}})$, and use roulette wheel selection [Lipowski and Lipowska, 2012] to choose parents. More specifically, the probability of selecting $S_{x_{i}}$ is:
@@ D.1 Experimental setting
We sampled 1,000 data points from each dataset for our experiments. No preprocessing was required prior to sampling, except for MMLU, due to its heterogeneous and often non-standard question formats, which are incompatible with our perturbation-based framework. We excluded items that lacked self-contained semantic meaning and thus were unsuitable for perturbation (e.g., “Which of the following statements is true?”), purely mathematical expressions without contextual meaning (e.g., “If $A=(1,2,3,4)$, let $B=\{(1,2),(1,3),(4,2)\}$. Then $B$ is”), or other irregular or ambiguous phrasing. Only questions with clearly worded, self-contained statements were retained.

### v2

@@ 3 Experiments
Our framework requires three inputs: $\psi$, $\mathcal{P}er$, and $a_{\text{Similarity}}$. For $\mathcal{P}er$, we apply two strategies: (1) ChatGPT-based paraphrasing, generating nine perturbed versions per question, (2) GAAP introduced in Sec. 2.6 with a similarity threshold of $\delta=0.7$. For $\psi$, we employ three state-of-the-art approaches: (i) SBERT-small (paraphrase-MiniLM-L6-v2), (ii) SBERT-large (all-mpnet-base-v2) (reimers2019sentence), and (iii) DeBERTa-v2-xlarge-MNLI (he2020DeBERTa). For (i) and (ii), we use cosine similarity $a_{\text{Similarity}}(x,x^{\prime})=(1+\cos(x,x^{\prime}))/2$. While, (iii) generates an entailment score; however, this score is not symmetric. To address this, we take the average $(a_{\text{Similarity}}(x,x^{\prime})+a_{\text{Similarity}}(x^{\prime},x))/2$. All experiments were conducted with an NVIDIA A100 GPU. Detailed experimental set up including prompts and parameters used are detailed in Appendix D. Appendix D also includes additional simulation results; we present only the core findings in the main paper for clarity and focus.
Table 2: Comparison of TSU across different temperature ranges for TriviaQA and MMLU.
Method TriviaQA MMLU
TSU(1.0,1.4) TSU(0.7,1.4) TSU(0.7–1.4) TSU(0.3–1.4) TSU(1.0,1.4) TSU(0.7,1.4) TSU(0.7–1.4) TSU(0.3–1.4)
Semantic Entropy 17.35 20.64 5.35 3.94 33.20 39.80 4.93 2.09
Kernel Entropy 43.92 55.56 18.23 9.64 59.48 69.10 37.32 12.63
VU 38.78 42.86 4.92 0.00 37.62 38.73 2.59 0.00
P(True) 3.85 3.49 0.00 0.00 5.87 5.79 0.00 0.00
LexSim 46.94 53.54 12.38 8.16 55.06 61.22 30.61 15.28
DegMat 45.37 47.96 20.02 13.27 69.39 77.55 32.58 14.34
LUQ 48.06 50.00 27.63 10.20 61.22 62.24 27.55 10.80
KLE 13.45 6.42 1.31 0.00 26.53 12.23 2.67 0.00
Inv-Entropy 77.55 88.78 47.21 19.05 73.47 86.73 43.88 18.37
NI-Entropy 61.22 60.20 19.32 11.73 50.00 59.38 18.37 7.14
WD-px-py 57.62 69.39 36.73 12.06 66.38 69.39 22.54 11.22
MAX-py-x 74.49 81.63 32.58 16.03 65.41 80.61 25.51 14.42
@@ 3.1 Results
Correctness-based
As shown in Table 1, Inv-Entropy achieves consistently strong and stable performance across all five datasets and all three metrics. It attains state-of-the-art results in both AUROC and PRR. The improvement is particularly clear on MMLU, where Inv-Entropy reaches 0.780 AUROC and 0.898 PRR, noticeably higher than all baselines, reflecting the advantage of our probabilistic framework with inverse design when the output information is limited. It also achieves leading performance on the long-answer dataset NQ, indicating that its effectiveness is not constrained by answer length (a detailed sensitivity analysis with respect to answer length is provided in Table 6 of the Appendix). It also ranks among the top two in Brier score, indicating well-calibrated confidence estimates. We intentionally use ChatGPT-based paraphrasing to highlight the advantages of our method independent of GAAP.
@@
Table 2 reports TSU results TriviaQA and MMLU. These results align with the correctness-based UQ metrics in Table 1: methods defined by our probabilistic framework consistently achieve top performance, while LexSim, DegMat and LUQ exhibit highly inconsistent results, occasionally leading in specific cases but underperforming in others. Methods like P(True) perform poorly in TSU due to their binary nature, which fundamentally lacks granularity. Similarly, Semantic Entropy shows limited discriminatory power as its values often cluster around few discrete points. In contrast to these coarse-grained approaches, Inv-Entropy is fine-grained with superior ability to reflect uncertainty variations, demonstrated by consistent top TSU performance across all datasets and temperature settings. This result highlights the probabilistic framework’s ability to capture uncertainty trends beyond labeled datasets. The complete TSU results are presented in Table 5 in the Appendix.
@@
(i) Impact of replications: The comparison between NR-Inv-Entropy and Inv-Entropy in Tables 1 and 2 highlights the impact of replication and bootstrapping. Notably, NR-Inv-Entropy often performs competitively, showcasing the strength of our framework even without replication. This suggests that one can trade off a small loss in accuracy for fewer queries. Appendix D includes an ablation over $S$ and $r$, further confirming this finding. (ii) Impact of inversion: The comparison between NI-Entropy and Inv-Entropy further confirms the advantages of our inverse approach, which examines the diversity of inputs that could have led to a specific output. Although NI-Entropy consistently underperforms Inv-Entropy, it still often ranks second, highlighting the strength of our defined probabilistic framework even without the inversion. (iii) Framework generality: The often competitive performance of WD-px-py and MAX-py-x highlights the robustness of our framework and its flexibility in defining a wide range of UQ measures. (iv) Impact of perturbations: Fig. 5 shows that GAAP consistently improves AUROC across all three datasets and both LLMs compared to ChatGPT-based paraphrasing. By enhancing input diversity in a principled way, GAAP better tests model robustness and improves uncertainty quantification. These results underscore the importance of meaningful perturbations. (v) Impact of embedding function: Our framework delivers consistently strong uncertainty estimates with every encoder we tested, including SBERT (paraphrase-MiniLM-L6-v2), SBERT (all-mpnet-base-v2), and DeBERTa. Although the larger encoders provide slightly higher scores on several tasks, the overall gap is modest, showing that even lightweight models can support reliable UQ when paired with our method. The close alignment between results from SBERT (all-mpnet-base-v2) and DeBERTa further suggests that entailment and similarity signals extracted by the two architectures contain overlapping information.
@@ 4 Conclusion
Appendix Outline
  • Appendix A: Related Work
  • Appendix B: Proof of Lemma 2.1
  • Appendix C: GAAP 2.6 Implementation Details
  • Appendix D: Settings & Additional Experimental Results:
    • Experimental setting D.1
    • Computational cost comparison D.2
    • Additional experimental results D.3
        Table 5: Complete TSU results.
        Table 6: Sensitivity Analysis to answer lengths in NQ.
        Table 7: Abalation study on the number of bootstrapping iterations S.
        Table 8: Abalation study on the number of replications r.
        Table 9: Abalation study on the embedding functions $\psi$.
        Table 10: Abalation study on temperatures $\mathbb{T}$ on TriviaQA.
        Table 11: Abalation study on temperatures $\mathbb{T}$ on SciQ.
@@ Appendix A Related Work
UQ in LLMs has been gaining increasing interest recently [huang2024survey]. The existing, albeit limited, literature can generally be categorized into three perspectives.
@@
Self-evaluation techniques for LLMs [chen2024] and the verbal expression of uncertainty [Lin2022] have recently been explored to enhance interpretability and reliability. For instance, recent approaches for evaluating free-form generation tasks frequently utilize LLMs as evaluators [arena-judge]. Additionally, for gray-box models, where the internal workings are partially known, perplexity [Chen1998] and entropy [shannon1948mathematical] can be directly computed from the output logits, providing a natural UQ measure.
@@
These methods generate multiple outputs for a given input and measure the deviation between them to estimate uncertainty [grewal2024, wagner2024, Kuhn2023]. Perhaps most prevalent is semantic entropy [farquhar2024detecting], which computes uncertainty by clustering semantically equivalent answers from multiple responses and calculating the entropy of the resulting clusters. While effective in capturing aleatoric uncertainty, these approaches struggle with confidently wrong predictions, as resampling often yields similar incorrect results, leading to overconfidence and poor calibration. This issue exacerbates the challenges of handling hallucinations in LLMs [grewal2024].
@@
This is a more recent approach that involves systematic perturbations of inputs or latent representations to evaluate output variability [zhang2023sac]. While dong2023revisit systematically evaluates LLM robustness for noisy slot-filling tasks under diverse input perturbations, SPUQ [gao2024spuq] provides a UQ metric by analyzing response variations to perturbed inputs. Notably, SPUQ achieves significant improvements in model uncertainty calibration, reducing Expected Calibration Error (ECE) by an average of 50%. Indeed, our theoretical argument in Sec. 2.1 provides a theoretical justification for the need for perturbation in effective UQ.
@@
By assumption, there exists a constant $G>0$ such that $\lVert\nabla^{2}\hat{f}(x)\rVert_{op}\leq G$, $\forall x\in\mathbb{R}^{d}$. Therefore,
@@
Since $z$ admits a $d$-dimensional isotropic Gaussian distribution $\mathcal{N}(0,\sigma^{2})$, we can provide an upper bound for $\mathbb{E}(R(z))$ and $\mathbb{E}(R^{2}(z))$ as
@@
$=\mathbb{E}\left[\left(\nabla\hat{f}(x_{0})^{\top}(I-\eta_{\nabla f^{\star}})z\right)^{2}\right]+2\mathbb{E}\left[\left(\nabla\hat{f}(x_{0})^{\top}(I-\eta_{\nabla f^{\star}})z\right)\zeta(z)\right]+\mathbb{E}\left[\zeta(z)^{2}\right],$
@@ C.1 An example
Table 3: A question $S_{x_{0}}$ and 1010 perturbed versions of $S_{x_{0}}$ as the output of GAAP.
@@ C.2.1 Key words selection
where the indices $\{j_{1},j_{2},\dots,j_{q}\}\subseteq\{1,\dots,p\}$, and the number of selected key tokens are: $q=\text{int}[r\cdot p]$. In GAAP, we use KeyBERT [keybert] to implement $k(\cdot,\cdot)$.
@@ C.2.3 Iterative population update
    In the terminology of genetic algorithms, we define our fitness function as the semantic similarity to $x_{0}$, $a_{\text{Similarity}}(x_{0},\cdot)$. Then, we compute the fitness value $a_{\text{Similarity}}(x_{0},x_{i})$ for $\forall S_{x_{i}}\in\text{Pop}_{t}(S_{x_{0}})$, and use roulette wheel selection [roulette] to choose parents. More specifically, the probability of selecting $S_{x_{i}}$ is:
@@ D.1 Experimental setting
No preprocessing was required except for MMLU, due to its heterogeneous and often non-standard question formats, which are incompatible with our perturbation-based framework. We excluded items that lacked self-contained semantic meaning and thus were unsuitable for perturbation (e.g., “Which of the following statements is true?”), purely mathematical expressions without contextual meaning (e.g., “If $A=(1,2,3,4)$, let $B=\{(1,2),(1,3),(4,2)\}$. Then $B$ is”), or other irregular or ambiguous phrasing. Only questions with clearly worded, self-contained statements were retained.